OpenStudy (anonymous):

I need help determining the linear factors and how many zero's there are in this problem... P(x)=X^45+x^44

OpenStudy (anonymous):

start with \[P(x)=x^{44}(x+1)\]

OpenStudy (anonymous):

where did the (x+1) come from??

OpenStudy (anonymous):

i was factoring.

OpenStudy (anonymous):

\[x^{44}(x+1)=x^{45}+x^{44}\]

OpenStudy (anonymous):

Can you show me how to use the descartes rule of signs..... that's what we've been using.... though I don't understand it

OpenStudy (anonymous):

so there are two zeros, namely at x = 0 and at x = -1

OpenStudy (anonymous):

don't need that here. just factor

OpenStudy (anonymous):

can you show me how? like draw it out?

OpenStudy (anonymous):

i am not sure what you mean. you have a polynomial with two terms, \[x^{45}\] and \[x^{44}\] and they have a common factor of \[x^{44}\]so you can "factor it out" just like you can write \[x^2+x=x(x+1)\]

OpenStudy (anonymous):

okay, so how many linear factors are there?